Calculate the potential of a cell where the Al3+ concentration is only 0.10 M. 

2 Al (s) + 3 Cu2+ (aq, 1.0 M) → 2 Al3+ (aq, 1.0 M) + 3 Cu (s)


1
Expert's answer
2022-03-25T05:23:03-0400

2 Al (s) + 3 Cu2+ (aq, 1.0 M) → 2 Al3+ (aq, 1.0 M) + 3 Cu (s)

Al(s) -3e = Al3+(aq), E0ox;

Cu2+(aq) +2e = Cu(s), E0red;

E0cell = E0red + E0ox;

Ecell = E0cell + R *T/(n * F) * lg(a2(Al3+)/a3(Cu2+)).

For T = 25 C = 298 K:

R *T/ F = 0.059

Ecell = E0cell + 0.059/3 * lg(0.12/13) = E0cell + 0,020*lg(0.01) = E0cell - 0.04 V.